FOR EXAMPLE:

C3H6O + 4O2 → 3CO2 + 3H2O ∆H=-1790kJ


Using 1 mole of C3H6O: -1790kJ

Using 2 moles of C3H6O: 2(-1790kJ)


Consider the information above on the combustion of acetone. Solve

the following problem involving the Enthalpy of Chemical Reactions. Show yours solutions. You may refer to the lecture notes for sample problems.


Calculate the amount of heat released during the combustion of

a. 0.75 mole of C3H6O

b. 120g of C3H6O


1
Expert's answer
2022-03-31T00:24:04-0400

a. If 1 mole of C3H6O: -1790kJ

(075/ 1)× -1790= -1342.5 kJ

b. 120g of C3H6O

Moles= 120/58= 2.07 moles

Heat = 2.07 × -1790= - 3703.45 kJ